# Chia-Hsiang Lin, , and Jhao-Ting Lin

### v1

@@ II-B Geometry-Based Algorithm Design
II-C Quantum Prism Design
Let us specify how we update $f$ in (II-B), whose architecture design and training strategy will be discussed in this section. If we could have a pretrained $f$, then the alternating step (II-B) could actually be omitted. However, the mission of $f$ is to perform the virtual light-splitting function, which does not associate with a real HSI sensor; that said, big data collection for pretraining $f$ does not seem to be feasible, due to the lack of ground-truth real HSI. This motivates us to train $f$ using the deep image prior (DIP) strategy [37], because this can be done by using just the single data (i.e., the observable MSI ${\bm{Z}}_{m}$). Another great advantage is that DIP is an unsupervised strategy, which can be done in a blind manner.
TABLE I: Quantum gates used in the quantum prism $f$, their symbols, and the corresponding unitary operators. In the mathematical definitions, let $\delta\triangleq\cos(\theta/2)$ and $\gamma\triangleq\sin(\theta/2)$.
@@ II-C Quantum Prism Design
To be more specific, the network $f$ accepts the input ${\bm{Z}}_{m}$, and outputs the virtual HSI ${\bm{Z}}_{h}$, merely from which we can unsupervisedly train/optimize the network parameters with the loss function provided by (II-B). Note that through the training/optimization of (II-B), the single data pair $({\bm{Z}}_{m},{\bm{Z}}_{h}^{t})$ is sufficient to complete the learning of the light-splitting relation. Since both the training data (i.e., ${\bm{Z}}_{m}$) and network architecture $f$ (to be designed next) are fixed, we do not need to train $f^{t+1}$ from scratch. Instead, we can initialize the network parameters as $f^{t}$ when training $f^{t+1}$ during the $(t+1)$th iteration of Algorithm 1. In other words, there are 10 iterations in PRIME, where $f$ is carefully trained with 100 epochs during the first iteration, while $f$ is fine-tuned with only 30 epochs in the remaining nine iterations. As the PRIME algorithm evolves as the iteration number $t$ increases, the virtual HSI ${\bm{Z}}_{h}^{t}$ gets better, and through the criterion (II-B), this better information of ${\bm{Z}}_{h}^{t}$ is injected into the learning of $f$, thereby obtaining a better light-splitting function $f^{t+1}$. In this work, (II-B) is optimized via the adaptive moment estimation (Adam) optimizer (learning rate set as $0.005$) [33].
@@
Figure 2: The illustrating architecture of the QUEEN-based [23] quantum prism network $f$, where the learning parameters include $(\rho_{i},\omega_{i},\theta_{i},\phi_{i})$. Note that $x_{i}$’s are not learning parameters but for data encoding, meaning that the feature value $x_{i}$ is encoded as the angle of the quantum rotation gate $R_{Y}$ (a.k.a. quantum angle embedding). DC module highly compresses the input image information into very few qubits, which are then fed into the core FE module with the cross-feature ordering specified in the yellow cubes. Provably, the core FE module can realize any valid quantum operators. Once the core quantum signal processings are done with the core FE module, we use the inverse-QC module to read out the well processed quantum signals (cf. [23, Section II.C]). Finally, the SS module is designed according to the aforementioned light-splitting relation of $0.5([{\bm{Z}}_{m}]_{(i,:)}\pm[{\bm{\theta}}]_{(i,:)})$ (cf. Section II-B) for shaping the spectra.
@@
Next, we design the QUEEN-based prism function $f$. We will adopt the quantum feature extraction technique developed in HyperQUEEN [23], partly because we believe that the advanced concept of quantum features can contribute to the learning of the complicated light-splitting task (a.k.a. spectral super-resolution), and partly because the low-rank module in HyperQUEEN may be interpreted as the spectral super-resolution module (cf. [23, Table II]). To get a sense, QUEEN is a deep network composed of quantum neurons, as its name suggested. Here, quantum neurons are also known as (a.k.a.) quantum logic gates. Some frequently used ones are summarized in Table II-C, including the rotation gates, Ising gate, Pauli gate, and Toffoli gate, etc. One can see that all these quantum gates correspond to some unitary operators [23, Section II], defined in Table II-C, where we also display the symbols of these gates.
@@
The design of the quantum prism network $f$ is graphically illustrated in Figure 2, which is composed of four modules, including the deep compression (DC) module, the core full expressibility (FE) module, the spectrum shaping (SS) module, and the inverse-QC module for addressing the quantum collapse (QC) effect, to be introduced below. The DC module is proposed to account for the very limited quantum bit (qubit) resources in current quantum computers; for example, the almost most advanced quantum computer called “IBM Osprey” has only 433 qubits. Thus, the DC module (defined in Table II) highly compresses the image information into very few qubits, for conducting the core quantum image processing in the highly compressed feature space [23]. The features $x_{i}$’s are then fed into the core FE module as the angles of the quantum rotation gates $R_{Y}$ (a.k.a. quantum angle embedding [38]) by using the cross-feature ordering specified in the yellow cubes of Figure 2. This cross-feature strategy is to allow more different features to interact with each other, thereby having a deeper understanding of related information [23]. The core FE module is more complicated, so we introduce it later.
@@
Once the core quantum signal processings are done with the core FE module, we use the inverse-QC module (defined in Table II) to read out the well processed quantum signals; the related theory can be found in [23, Section II.C]. Simply speaking, the QC effect says that once we attempt to read/observe the quantum state of the processed signals, it will immediately collapse to some eigenstate; thus, the function of the inverse-QC module aims at learning the inverse mapping from some measuring statistics (associated with the collapsed quantum state) to the target image [23]. On the other hand, the SS module (defined in Table II) is designed according to the aforementioned light-splitting relation that splits an MSI band into $\gamma:=2$ HSI bands, i.e., $0.5([{\bm{Z}}_{m}]_{(i,:)}\pm[{\bm{\theta}}]_{(i,:)})$ (cf. Section II-B). This relation allows us to much more efficiently focus on learning only half of the bands (say, even bands), followed by generating the remaining odd bands for alternately inserting into the even bands, thereby generating the virtual HSI ${\bm{Z}}_{h}=f({\bm{Z}}_{m})$, as graphically illustrated in the lower-right block of Figure 2.
@@
To well process the quantum signals, we need a strong core processing module, such as a module with FE [23, Theorem 2]. Therefore, we employ a recently developed Ising-Rotation quantum network architecture to design our core quantum FE module, sequentially composed of Rotation Y gate, Ising XX gate, Rotation X gate, Ising XX gate, and Rotation Y gate. We then have the mathematical guarantee that the core FE module can express any valid quantum operators, as stated in the following theorem.
@@
The proof exactly follows the procedure of proving [23, Theorem 2], and is omitted here for conciseness. As also proved in [23], the Toffoli quantum entanglement is useful, and hence utilized in our core FE module as well. In the last stage, the Pauli-Z gate is used to measure the quantum information using a strategy similar to the max pooling, and such a strategy is also proven useful even in some rather complicated inverse imaging tasks [23]. Overall, the core FE module is defined in Table II. Therefore, we have completed the design of the QUEEN-based quantum prism network $f$.
@@
The interested readers are referred to [23, Section II.A] and [23, Section II.B], where a comprehensive review of the fundamental concepts required in QUEEN is presented, including quantum deep learning, barren plateaus effect, Dirac notation system, and quantum measurement/collapse, etc. Next, we experimentally demonstrate the strength of this QUEEN-based PRIME algorithm for MU.
TABLE II: Detailed architecture of the proposed quantum prism $f$. “CB($c$,$s$,$p$)” denotes convolution block composed of the convolution layer “Conv($c$,$s$,$p$)” and LeakyReLU (with negative slope 0.2), while “TCB($c$,$s$)” denotes transposed CB composed of transposed convolution TConv($c$,$s$) and LeakyReLU, where $(c,s,p)$ specify the number of output channels, the kernel size of $s\times s$, and the number of paddings, respectively. The “MaxPool” denotes the Max Pooling, the non-negative projection (“NNProj”) is implemented by ReLU, and the “Up-sample” is implemented by bilinear interpolation, respectively.

### v2

@@ III-B Geometry-Based Algorithm Design
III-C Quantum Prism Design
Let us specify how we update $f$ in (III-B), whose architecture design and training strategy will be discussed in this section. If we could have a pretrained $f$, then the alternating step (III-B) could actually be omitted. However, the mission of $f$ is to perform the virtual light-splitting function, which does not associate with a real HSI sensor; that said, big data collection for pretraining $f$ does not seem to be feasible, due to the lack of ground-truth real HSI. This motivates us to train $f$ using the deep image prior (DIP) strategy [42], because this can be done by using just the single data (i.e., the observable MSI ${\bm{Z}}_{m}$). Another great advantage is that DIP is an unsupervised strategy, which can be done in a blind manner.
TABLE I: Quantum gates used in the quantum prism $f$, their symbols, and the corresponding unitary operators. In the mathematical definitions, let $\delta\triangleq\cos(\theta/2)$ and $\zeta\triangleq\sin(\theta/2)$.
@@ III-C Quantum Prism Design
To be more specific, the network $f$ accepts the input ${\bm{Z}}_{m}$, and outputs the virtual HSI ${\bm{Z}}_{h}$, merely from which we can unsupervisedly train/optimize the network parameters with the loss function provided by (III-B). Note that through the training/optimization of (III-B), the single data pair $({\bm{Z}}_{m},{\bm{Z}}_{h}^{t})$ is sufficient to complete the learning of the light-splitting relation. Since both the training data (i.e., ${\bm{Z}}_{m}$) and network architecture $f$ (to be designed next) are fixed, we do not need to train $f^{t+1}$ from scratch. Instead, we can initialize the network parameters as $f^{t}$ when training $f^{t+1}$ during the $(t+1)$th iteration of Algorithm 1. In other words, there are $10$ iterations in PRIME, where $f$ is carefully trained with $100$ epochs during the first iteration, while $f$ is fine-tuned with only $30$ epochs in the remaining nine iterations. As the PRIME algorithm evolves as the iteration number $t$ increases, the virtual HSI ${\bm{Z}}_{h}^{t}$ gets better, and through the criterion (III-B), this better information of ${\bm{Z}}_{h}^{t}$ is injected into the learning of $f$, thereby obtaining a better light-splitting function $f^{t+1}$. In this work, (III-B) is optimized via the Adam optimizer (learning rate set as $0.005$) [38].
@@
Figure 2: The illustrating architecture of the QUEEN-based [8] quantum prism network $f$, where the learning parameters include $(\rho_{i},\omega_{i},\theta_{i},\phi_{i})$. Note that $x_{i}$’s are not learning parameters but for data encoding, meaning that the feature value $x_{i}$ is encoded as the angle of the quantum rotation gate $R_{Y}$ (i.e., quantum angle embedding). DC module highly compresses the input image information into very few qubits, which are then fed into the core FE module with the cross-feature ordering specified in the yellow cubes. Provably, the core FE module can realize any valid quantum operators. Once the core quantum signal processings are done with the core FE module, we use the inverse-QC module to read out the well processed quantum signals (cf. [8, Section II-C]). Finally, the SS module is designed according to the light-splitting equation of $0.5([{\bm{Z}}_{m}]_{(i,:)}\pm[{\bm{\theta}}]_{(i,:)})$ (detailed in Appendix -A) for shaping the pixel spectra.
@@
Next, we design the QUEEN-based prism function $f$. We will adopt the quantum feature extraction technique developed in HyperQUEEN [8], partly because we believe that the advanced concept of quantum features can contribute to the learning of the complicated light-splitting task (i.e., spectral super-resolution), and partly because the low-rank module in HyperQUEEN may be interpreted as the spectral super-resolution module (cf. [8, Table II]). To get a sense, QUEEN is a deep network composed of quantum neurons, as its name suggested. Here, quantum neurons are commonly referred to as quantum logic gates. The critical ones used in $f$ are summarized in Table III-C, including the rotation gates, Ising gate, Pauli gate, and Toffoli gate, etc. One can see that all these quantum gates correspond to some unitary operators [8, Section II], defined in Table III-C, where we also display the symbols of these gates.
@@
The design of the quantum prism network $f$ is graphically illustrated in Figure 2, which is composed of four modules, including the deep compression (DC) module, the core full expressibility (FE) module, the spectrum shaping (SS) module, and the inverse-QC module for addressing the quantum collapse (QC) effect. To explicitly describe the network architecture for reproducibility while maintaining the readability of the paper, we provide the network details (e.g., configuration and output size of each layer) in Appendix -B. The DC module is proposed to account for the very limited quantum bit (qubit) resources in current quantum computers; for example, the almost most advanced quantum computer called “IBM Osprey” has only 433 qubits. Thus, the DC module (defined in Appendix -B) highly compresses the image information into very few qubits, for conducting the core quantum image processing in the highly compressed feature space [8]. The features $x_{i}$’s are then fed into the core FE module as the angles of the quantum rotation gates $R_{Y}$ (i.e., quantum angle embedding [43]) by using the cross-feature ordering specified in the yellow cubes of Figure 2. This cross-feature strategy is to allow more different features to interact with each other, thereby having a deeper understanding of related information [8]. The core FE module is more complicated, so we introduce it later.
@@
Once the core quantum signal processings are done with the core FE module, we use the inverse-QC module (defined in Appendix -B) to read out the well processed quantum signals; the related theory can be found in [8, Section II-C]. Simply speaking, the QC effect says that once we attempt to read/observe the quantum state of the processed signals, it will immediately collapse to some eigenstate; thus, the function of the inverse-QC module aims at learning the inverse mapping from some measuring statistics (associated with the collapsed quantum state) to the target image [8]. On the other hand, the SS module (defined in Appendix -B) is designed according to the aforementioned light-splitting relation that splits an MSI band into $\gamma:=2$ HSI bands ($\gamma$ is the upsampling factor in the spectral super-resolution task), i.e., $0.5([{\bm{Z}}_{m}]_{(i,:)}\pm[{\bm{\theta}}]_{(i,:)})$ (cf. Appendix -A). This relation allows us to much more efficiently focus on learning only half of the bands (e.g., the odd bands), followed by generating the remaining even bands for alternately inserting in between the odd bands, thereby generating the complete virtual HSI ${\bm{Z}}_{h}=f({\bm{Z}}_{m})$, as graphically illustrated in the lower-right block of Figure 2.
@@
To well process the quantum signals, we need a strong core processing module, such as a module with FE [8, Theorem 2]. Therefore, we employ a recently developed Ising-Rotation quantum network architecture to design our core quantum FE module, sequentially composed of Rotation Y gate, Ising XX gate, Rotation X gate, Ising XX gate, and Rotation Y gate. We then have the mathematical guarantee that the core FE module can express any valid quantum operators, as stated in the following theorem.
@@
The proof exactly follows the procedure of proving [8, Theorem 2], and is omitted here for conciseness. As also proved in [8], the Toffoli quantum entanglement is useful, and hence utilized in our core FE module as well. In the last stage, the Pauli-Z gate is used to measure the quantum information using a strategy similar to the max pooling, and such a strategy is also proven useful even in some rather complicated inverse imaging tasks [8]. Overall, the core FE module is defined in Appendix -B. Therefore, we have completed the design of the QUEEN-based quantum prism network $f$.
@@
The interested readers are referred to [8, Section II-A] and [8, Section II-B], where a comprehensive review of the fundamental concepts required in QUEEN is presented, including quantum deep learning, barren plateaus effect, Dirac notation system, and quantum measurement/collapse, etc. Next, we experimentally demonstrate the strength of this QUEEN-based PRIME algorithm for MU.
IV Experimental Results
This section aims to demonstrate the effectiveness of the proposed PRIME algorithm. As there is no MU work targeted at the underdetermined case, we propose an experimental protocol for evaluating the performance of the MU algorithms in Section IV-A, where one can also find the detailed experimental setting and data description. In Section IV-B, we conduct comprehensive comparisons with the MU methods used in [3]. The ablation study for the PRIME algorithm is investigated in Section IV-C. Finally, in Section IV-D, we show that the proposed PRIME algorithm is generally applicable to different sensors, and analyze the impact of the upsampling factor $\gamma$ on the performance of PRIME.
IV-A Experimental Protocol and Design
Since prior MU works do not consider the underdetermined scenario, we cannot find suitable experimental protocols to evaluate the MU performances. In view of this, we design a protocol (cf. Figure 3) to facilitate future development of the critical MU technique. Since there is no benchmark MU dataset with available ground-truth (GT) ${\bm{B}}$ and ${\bm{S}}$, our protocol (referred to as Lin’s protocol) is particularly designed to address this issue. This protocol is inspired by Wald’s protocol, which has been widely used for MSI/HSI pansharpening quality evaluation [44, 45, 46, 47]. Specifically, we adopt the reduced resolution (RR) assessment [46, Figure 2], which mirrors the synthesis property of Wald’s protocol. Specifically, the process begins with a reference HSI, from which effective HU methods (e.g., HISUN) are used to first generate the GT abundances ${\bm{S}}$ and reference hyperspectral signatures ${\bm{A}}_{\textrm{ref}}$, where ${\bm{A}}_{\textrm{ref}}$ is further uniformly downsampled along the spectral dimension to obtain the GT multispectral signatures (endmembers) in ${\bm{B}}$. Next, ${\bm{S}}$ and ${\bm{B}}$ are used to synthesize the reference MSI ${\bm{Z}}_{m}={\bm{B}}{\bm{S}}$, which are then fed into the MU methods to estimate the signatures $\widehat{{\bm{B}}}$ and abundances $\widehat{{\bm{S}}}$. Finally, the similarity between $\widehat{{\bm{S}}}$ and ${{\bm{S}}}$, as well as the similarity between $\widehat{{\bm{B}}}$ and ${{\bm{B}}}$, are measured with suitable quality metrics, as illustrated in Figure 3.
For the reference HSI to be used in the protocol, we consider the data acquired by NASA’s Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) sensor [48]. The AVIRIS sensor is composed of $224$ contiguous bands, and only $172$ bands are used in our experiments after removing those water-vapor absorption bands (i.e., bands 1-10, 104-116, 152-170, and 215-224). Then, the spectral downsampling stage in the protocol is implemented using a uniform spectral downsampling matrix, which is widely accepted in the remote sensing area [9, 49]. This approach effectively and fairly merges spectral information across specified wavelength ranges to generate representative bands. Precisely, the spectral ranges correspond to the Landsat Thematic Mapper (TM) bands 1–4, covering 450–520, 520–600, 630–690, and 760–900 nm, respectively, are used to uniformly downsample the reference signatures ${\bm{A}}_{\text{ref}}$, resulting in the $4$-band GT multispectral signatures ${\bm{B}}$. Three data from diverse regions are introduced below. The first data (with a spatial resolution of 17.3 m) was acquired over Vancouver Island, British Columbia, Canada. The second data (with a spatial resolution of 11.7 m) was acquired over Montrose, CO, USA. The third data (with a spatial resolution of 16.8 m) was acquired over Nassau, FL, USA. All the three reference images are with $L=256\times 256$ pixels, and their false-color images (composed of band 20 (R), band 12 (G), and band 6 (B)) are displayed in Figure 4.
IV-B Qualitative and Quantitative Analysis
In this section, we conduct qualitative and quantitative analysis for the proposed PRIME algorithm (i.e., Algorithm 1) together with related methods, on the three $4$-band MSI data described in Section IV-A (i.e., $P:=4$). Also, we empirically set $N:=6$ for all the three data. There are rather few existing MU methods. We will compare with the MU methods used in [3], which extends VCA [17] and NMF [51] to conduct the MU task. Note that VCA will first perform dimensionality reduction (DR). Considering that the number of bands is smaller than the sources in MU (i.e., $P<N$), we ignore this DR step by running the algorithm in the $P$-dimensional space. Since VCA can only compute the signatures ${\bm{B}}$, we compute the abundance ${\bm{S}}$ by left-multiplying ${\bm{Z}}_{m}$ with the pseudo-inverse of ${\bm{B}}$, as suggested by the original paper [17]. As for NMF, the iteration number is set to $1000$ for better convergence. Besides, ${\bm{B}}$ and ${\bm{S}}$ in NMF are initialized by the powerful HISUN as a warm start strategy [1]. As NMF itself cannot yield a good abundance estimation result, we further refine ${\bm{S}}$ by solving the non-negative least-squares problem using the optimized ${\bm{B}}$ returned from NMF, in order to upgrade the performance of NMF.
With the judicious prism mechanism (cf. Figure 1), the induced MU data-fitting term (III-A) can well tackle the highly challenging underdetermined MU problem. Specifically, for all the three data, it can be seen that the proposed PRIME algorithm achieves highly promising results on all the signature/abundance estimations. Though the 3rd estimated signature in the Montrose data has some deviation, its curve shape still holds a high resemblance to the GT signature; see the red and black curves in Figure 6(c). Accordingly, the 3rd abundance map estimated by PRIME still has very high quality, as can be seen from the first map (GT) and the second map (PRIME) in Figure 6(c). The highly remarkable thing is that the proposed PRIME algorithm successfully separates $6$ sources merely from $4$ observations (i.e., $N>P$), demonstrating the feasibility of MU for the underdetermined case, under a fully blind setting!
To fairly argue the superior performance of PRIME, we also conduct a quantitative evaluation. The estimated results over three data in terms of SAM, RMSE, and computational time are summarized in Table II, where the boldfaced numbers indicate the smallest SAM/RMSE or the fastest computational time. For all three data, one can observe that the proposed PRIME algorithm achieves the smallest SAM, indicating the best spectral shape preservation capability in $\widehat{{\bm{B}}}$, as well as the smallest RMSE, revealing the excellent spatial quality of the estimated abundances in $\widehat{{\bm{S}}}$. The computational time of PRIME, however, is longer than other fast algorithms like VCA and NMF, mainly owing to the iterative network update stage (cf. (III-B)). Nevertheless, if one accepts a non-blind mechanism, the quantum prism network $f$ may be pre-trained, and we expect that the overall computational time of PRIME would be reduced to just a few seconds.
TABLE II: Performance comparison among different methods in terms of SAM, RMSE, and computational time.
[…]
Vancouver Island VCA 13.3257 0.3027 0.0137
NMF 12.8905 0.2455 18.9011
PRIME 6.5405 0.1612 54.6156
Montrose VCA 3.8597 0.1628 0.0136
NMF 4.1394 0.2001 16.5596
PRIME 2.2104 0.0916 54.6125
Nassau VCA 13.1555 0.2940 0.0127
NMF 20.1224 0.4466 18.4992
PRIME 8.1568 0.0830 54.9062
IV-C Ablation Study
In this experiment, we conduct an ablation study to investigate the influence of three key components of PRIME. First, PRIME has adopted the heuristic initialization (HI) strategy (13). We will show that this is better than the random initialization strategy when computing ${\bm{Z}}_{h}^{0}$. Second, we evaluate the value of the SS module in the QUEEN-based prism $f$, which judiciously utilizes the physical meaning of light splitting to reduce half of the network parameters as described in Section III-C. Otherwise, when SS module is removed, the last layer of the inverse-QC module (i.e., TConvModule 3 defined in Appendix -B) should be revised to directly have $M=8$ output channels. Though this seems more straightforward, the performance significantly degrades, as will be seen. Third, to implement the tough subproblem (10), we update $({\bm{A}}^{t},{\bm{S}}^{t})$ by using the convex geometry (CG) theory [20]. Otherwise, we need to implement (10) directly using the algebra-based NMF method [51]. Nevertheless, as will be seen, this is less efficient.
We test the proposed PRIME algorithm by turning on two of the three components (and turning off the other one) for all the three data introduced in Section IV-A. The ablation study results averaged over the three data are quantitatively displayed in Table III. We observe that the MU performance of PRIME significantly degrades if SS module or CG module is turned down. Though the performance of PRIME also degrades when HI module is not used, it is not too significant. This shows that initialization does play some role when implementing the non-convex MU criterion (III-A), but even when ${\bm{Z}}_{h}^{0}$ is simply initialized by uniformly random matrix, the proposed PRIME algorithm still has the capability of retrieving relatively good estimates of $(\widehat{{\bm{B}}},\widehat{{\bm{S}}})$ through the subsequent iterative updating procedure (cf. Algorithm 1). The proposed PRIME algorithm achieves the best MU performance when all the HI/SS/CG modules are used, proving the effectiveness of these components, as can be seen in Table III.
TABLE III: Ablation study for the proposed PRIME algorithm, where the marker “✔” means that the corresponding module is used.
[…]
20.4113 0.5307
14.4840 1.5109
12.1527 0.1937
5.6359 0.1119
Figure 8: Region of interest (ROI) of the studied Pavia data acquired by the PRISMA satellite operated by the Agenzia Spaziale Italiana (ASI).
Figure 9: The MU results, including the estimated abundance maps and the multispectral endmember signatures, for the Pavia data (PRISMA satellite).
IV-D Related Discussions
To further evaluate the performance of PRIME, we use the reference HSI captured by PRecursore IperSpettrale della Missione Applicativa (PRISMA) [52] satellite, which was launched by the Agenzia Spaziale Italiana (ASI) in 2019. PRISMA is equipped with $66$ and $170$ contiguous spectral bands in the VNIR and SWIR spectral range, respectively. As well as Section IV-A, the spectral ranges corresponding to the Landsat TM bands 1–4 are used to uniformly downsample the reference signatures, resulting in the $4$-band GT multispectral signatures. The data (with a spatial resolution of 30 m) was acquired over Pavia, Italy, and with $L=256\times 256$ pixels, and the false-color image (composed of band 33 (R), band 21 (G), and band 12 (B)) is displayed in Figure 8.
The estimated multispectral signatures $\widehat{{\bm{B}}}$ and the corresponding abundance maps $\widehat{{\bm{S}}}$ are collectively displayed in Figure 9. For VCA, the 2nd signature entirely deviates from the GT (cf. Figures 9(b)). In addition, the 3rd, 5th, and 6th abundance maps appear to be highly similar to each other, lacking distinct spatial differences for distinguishing the materials/substances (cf. Figures 9(c), 9(e), and 9(f)). For NMF, the 1st, 2nd, and 5th signature estimations (along with their respective abundance maps) are fail, exhibiting notable discrepancies from the GT (cf. Figures 9(a), 9(b) and 9(e)). In contrast, it is evident that all the estimated signatures of PRIME closely resemble the GT. Moreover, PRIME demonstrates a strong ability to preserve spatial uniqueness across all abundance maps while holding high resemblance to GT. Table IV summarizes the quantitative evaluation results, revealing that PRIME outperforms peer methods by achieving the smallest SAM and RMSE. The evaluation using PRISMA satellite data further underscores the generalizability of PRIME, demonstrating its robustness across diverse sensing platforms.
TABLE IV: Performance comparison among different methods in terms of SAM, RMSE, and computational time.
[…]
Pavia VCA 17.2164 0.2067 0.0162
NMF 21.8513 0.1641 14.8462
PRIME 7.7992 0.1116 53.4070
Since we mainly focus on unmixing $N=6$ sources from $P=4$ band MSI in Section III, the $\gamma$ is set to $2$, which is sufficient to have $\gamma P>N$. However, some datasets, such as the Cuprite mining site collected by AVIRIS, may have $N\geq 10$ minerals [20], meaning that a larger value of $\gamma$ is needed. Therefore, we test PRIME on Montrose data to understand the impact of $\gamma$ by largely increasing its value from $2$ to $6$, as demonstrated in Figure 10. To accommodate the variation, we adjust the methodology accordingly, particularly within the quantum prism network $f$.
As the initial virtual HSI ${\bm{Z}}_{h}^{0}$ is relevant to the $f$ update (cf. Step 4 in Algorithm 1), we first modify the initialization process (cf. Step 2 in Algorithm 1) so that each multispectral band in ${\bm{Z}}_{m}$ is divided into $\gamma$ hyperspectral bands, while maintaining the light-splitting property with slight perturbation (cf. Appendix -A). Next, we revise the last layer of the inverse-QC module (i.e., TConvModule 3, as detailed in Appendix -B) to output ($\gamma P-P$) bands, while the remaining $P$ bands were calculated within the SS module based on the light-splitting relation. To maintain the spectral continuity of the virtual HSI, as $\gamma$ increased from $2$ to $6$ (cf. Figure 10), we progressively adjusted $\alpha$ to 1E-4 (for $\gamma=2$), 2E-4 (for $\gamma=3$), 5E-4 (for $\gamma=4$), 6E-4 (for $\gamma=5$) and 1E-3 (for $\gamma=6$). We observe that SAM and RMSE only have subtle variations as $\gamma$ increases, meaning that PRIME is able to address a more serious underdetermined BSS scenario that requires higher $\gamma$. In addition, the computational time approximately ranges from $54$ to $64$ seconds, indicating that the impact of $\gamma$ on the computational efficiency is ignorable. These findings highlight the robustness of the proposed PRIME algorithm, making it applicable for underdetermined MU scenarios requiring higher value of $\gamma$.
V Conclusion
In this work, we have the first attempt to resolve the MU problem for the underdetermined scenario, where we have more sources than the MSI bands. The proposed PRIME algorithm judiciously introduces a virtual quantum prism to tackle the underdetermined issue. The prism conducts the light-splitting task to generate more virtual bands, and induces a novel data-fitting term for MU. The data-fitting term, together with some customized geometry-based volume and algebra-based sparsity regularization terms, forms the proposed MU criterion. Our PRIME algorithm elegantly implements the criterion by plugging in the computationally efficient convex geometry method, and proposes a quantum deep image prior (QDIP) strategy to achieve a fully blind mechanism. For the virtual quantum prism, the adopted quantum feature extraction technique has the mathematical guarantee of quantum full expressibility. To evaluate the MU performance, we also design the first MU experimental protocol that can be easily implemented for future investigators. As it turns out, the PRIME algorithm is highly effective for addressing the MU problem in terms of both quantitative and qualitative perspectives.
Figure 10: Performance of PRIME tested across different settings of $\gamma$.
-A Initialization of PRIME Algorithm
Since the network $f$ will be updated first in Algorithm 1, we need to initialize ${\bm{Z}}_{h}$ and $({\bm{A}},{\bm{S}})$. For ${\bm{Z}}_{h}$, we propose to initialize it as a slightly perturbed light-splitting version of the $P$-band ${\bm{Z}}_{m}$, as detailed below. As mentioned in Section I, the virtual ${\bm{Z}}_{h}$ has no need to correspond to some real HSI sensor, so we can flexibly assume that each band in ${\bm{Z}}_{m}$ is split into $\gamma\in\mathbb{Z}_{++}$ bands (i.e., $M=\gamma P$, where $\gamma$ is equivalent to upsampling factor in spectral super-resolution task).
Therefore, the spectral response matrix ${\bm{D}}$ is given by ${\bm{D}}\triangleq{\bm{I}}_{P}\otimes\bm{1}^{T}_{\gamma}\in\mathbb{R}^{P\times M}$. Considering that the concept of the $M$-band virtual ${\bm{Z}}_{h}$ is proposed to mitigate the underdetermined issue (thus requiring $M=\gamma P\geq N$), and that we usually have $2P\geq N$ in practice as discussed before, we can simply consider $\gamma:=2$ for conciseness. According to the physical meaning of light splitting, the $i$th MSI band $[{\bm{Z}}_{m}]_{(i,:)}$ can be split into two HSI bands, i.e., $0.5([{\bm{Z}}_{m}]_{(i,:)}\pm{[\bm{\theta}]}_{(i,:)})$ for some $[\bm{\theta}]_{(i,:)}$, which ensures the light-splitting relation of ${\bm{D}}{\bm{Z}}_{h}={\bm{Z}}_{m}$. Considering the desired continuity of the HSI bands, we can set $[\bm{\theta}]_{(i,:)}$ as $\frac{1}{4}([{\bm{Z}}_{m}]_{(i+1,:)}-[{\bm{Z}}_{m}]_{(i,:)})$ for uniform spectral sampling; further considering the boundary issue, the last term (i.e., the $P$th term) is set as $[\bm{\theta}]_{(P,:)}$ as $\frac{1}{4}([{\bm{Z}}_{m}]_{(P,:)}-[{\bm{Z}}_{m}]_{(P-1,:)})$. To be precise, the above generated $M$-band information $\widetilde{{\bm{Z}}}_{h}$ can be explicitly written as
It can be proven true that the $M$-band $\widetilde{{\bm{Z}}}_{h}$ is still of rank $P$, so it is slightly perturbed with Gaussian noise ${\bm{N}}$ whose energy is about $e:=5\%$ of $\widetilde{{\bm{Z}}}_{h}$, leading to the initialization of
${\bm{Z}}_{h}^{0}=\widetilde{{\bm{Z}}}_{h}+c{\bm{N}},~{}\textrm{with}~{}c\triangleq\sqrt{\frac{e\|\widetilde{{\bm{Z}}}_{h}\|_{F}^{2}}{\|{\bm{N}}\|_{F}^{2}}},$ (13)
-B Detailed Design of Quantum Prism Network $f$
[…]
TABLE V: Detailed architecture of the proposed quantum prism $f$.